OpenStudy (anonymous):

If the Ellipse at the White House is 616 ft long and 528 ft wide, what is its equation? For simplicity's sake, model the center at (0,0) with the x-axis as its focal axis.

OpenStudy (alexwee123):

p.652

OpenStudy (anonymous):

That's not an equation

OpenStudy (alexwee123):

i know that exact problem is in p.652 in my book

OpenStudy (amistre64):

lol, well then its the correct answer :) but lets assume the book is a secret

OpenStudy (anonymous):

I don't have the book, just a paper from school.

OpenStudy (amistre64):

there are really 2 things we need to determine to define this equation; conventionally we call them a and b

OpenStudy (amistre64):



OpenStudy (anonymous):

x^2/(308^2) + y^2/(264^2)? Would that be it?

OpenStudy (amistre64):

with these the equation is defines as: \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\] where a is the longer and b the shorter values ... and they can change depending on if the ellipse is long up or long sideways

OpenStudy (anonymous):

Thank you

OpenStudy (amistre64):

so just split those values in half that they give for long and wide ... then square them

OpenStudy (amistre64):

good luck